# NaVIDA: Vision-Language Navigation with Inverse Dynamics Augmentation

### v1

@@ 4 Experiment
NaVILA [8] RSS2025 8B $\checkmark$ 6.77 49.3 44.0 58.8
UniNaVid [45] RSS2025 7B $\checkmark$ 6.24 48.7 40.9 -
Aux-Think [33] NeurIPS2025 8B $\checkmark$ 6.24 52.2 40.2 -
StreamVLN [39] arXiv2507 7B $\checkmark$ 6.22 52.9 46.0 61.9
MonoDream [34] arXiv2508 2B $\checkmark$ 6.38 49.4 40.9 -
InternVLA-N1(S2) [10] - 7B $\checkmark$ 6.41 49.5 41.8 62.6
InternVLA-N1(S1+S2) [10] - 8B $\checkmark$ $\checkmark$ 5.91 53.5 46.1 65.3
NavFoM [44] arXiv2509 7B $\checkmark$ 5.51 57.4 49.4 60.2
JanusVLN [43] arXiv2509 8B $\checkmark$ 6.06 56.2 47.5 62.1
@@ 4.1 Experiments Setup
Implementation Details. We implement NaVIDA on top of Qwen-2.5-VL-3B [4]. During training, the vision encoder is frozen, while the projector and LLM backbone are optimized with a learning rate of 2e52e-5 for one epoch over mixed VLN and inverse-dynamics data. Following NaVILA [8], we cap the temporal context at eight historical frames; if a sequence exceeds this limit, we uniformly subsample it to eight frames to preserve coverage across the full trajectory. Unless otherwise stated, HPAC uses merge probability $p=0.7$ and a maximum level-2 chunk size of $n=3$. At inference, we apply entropy-guided execution, which adaptively truncates the decoded chunk list to prevent long-horizon drift. Further hyperparameter details are provided in the Supplementary Material.
Evaluation Benchmarks. We evaluate NaVIDAon VLN-CE using the val-unseen splits of R2R-CE [3, 19] and RxR-CE [20, 19]. Following standard practice [43, 10, 33], we report Navigation Error (NE), Success Rate (SR), Oracle Success rate (OS), Success weighted by Path Length (SPL) and normalized Dynamic Time Warping (nDTW). These metrics jointly capture goal completion, trajectory efficiency, and path fidelity, and allow direct comparison to recent RGB-only and multi-sensor approaches. Detailed explanations of the metrics are provided in the Supplementary Material.
@@ 4.2 Main Results
Performance Evaluation. As shown in  Tab. 1 (R2R-CE) and  Tab. 2 (RxR-CE), NaVIDAdelivers state-of-the-art results on the val-unseen splits while using only a single RGB stream and a 3B backbone VLM. On R2R-CE, NaVIDAattains the best NE/OS/SR and the second-best SPL, surpassing the strongest RGB-only 8B baseline JanusVLN [43] by -0.46 NE, +4.3 OS, and +0.9 SR. These gains demonstrate that modeling action-vision causality improves goal attainment without sacrificing path efficiency. On RxR-CE, NaVIDA matches the best SR (57.4) and achieves the best NE/SPL/nDTW, improving over StreamVLN [39] and JanusVLN by +1.2 to +4.5 SR, +2.1 to +3.6 SPL, and -0.83 to -0.99 NE. The consistent advantage on the linguistically richer, longer-horizon RxR-CE indicates stronger long-range alignment and reduced drift, attributable to IDS and hierarchical chunking.
Compared with multi-sensor (e.g., ETPNav [2] and NavMorph [42]), NaVIDA’s single-RGB policy still leads on core metrics in both benchmarks, despite several competitors using panoramic RGB, depth, or odometry. Moreover, compared with other single RGB input methods, NaVIDA achieves the best results with  35% of the parameters and without auxiliary generators or spatial encoders. This further highlights the efficiency of action-grounded supervision as a training paradigm rather than a capacity increase. In summary, NaVIDA consistently outperforms strong RGB-only baselines and methods with richer sensors, establishing that inverse-dynamics augmentation plus HPAC forms an effective and lightweight alternative to conventional map/history cues and forward world modeling.
Table 3: Ablation study of IDS on VLN-CE R2R Val-Unseen split. We denote training without IDS by $\circ$, and with IDS by $\bullet$. Blank cells indicate the dataset is not used.
@@
Qualitative Results. Fig. 4 contrasts NaVIDA with a variant without IDS where green arrows mark correct actions and red arrows mark errors. NaVIDA correctly defers turning until clearing the staircase and selects the intended intersection, whereas the ablated model exhibits premature turns and junction misclassification. These are typical symptoms of error accumulation when future visual consequences of actions are not internalized.
Fig. 5 presents examples of real-world trials. NaVIDA executes tight turns at narrow corners and passes smoothly through single-person doorways; even with pedestrians ahead, it resists interference and holds a reliable heading. Taken together, the simulator and real-world evidence identify IDS as the key component behind the observed stability and scene generalization, outperforming the variant without inverse-dynamics supervision.
@@ 4.3 Ablation Study
In this section, we conduct ablations on the effectiveness of IDS, the length of HPAC, entropy-based execution horizon selection and dynamic modeling methods. To ensure efficiency, we conduct experiments on data without ScaleVLN and report the performance on R2R-CE benchmark. We do not use entropy-based execution horizon selection mechanism by default unless otherwise stated.
Effect of IDS Data.  Tab. 3 varies whether R2R, RxR, and DAgger trajectories contribute IDS pairs ($\bullet$) or not ($\circ$). Using IDS consistently improves all metrics: on R2R alone, SR rises from 32.0 to 40.5; with R2R+RxR, SR increases to 47.7; adding DAgger without IDS yields SR 52.6, while enabling IDS across all three reaches SR 55.5 with lower NE (4.91) and higher SPL (49.6). The performance gains are most pronounced in data-scarce scenarios and continue to accumulate as additional data sources are introduced. This pattern suggests IDS injects a transferable, language-agnostic notion of “how actions change what is seen,” which complements imitation rollouts such as DAgger.
Ablations of the Chunk Number. Table 4 shows the impact of the maximum number of action chunks allowed in HPAC on the VLN-CE R2R Val-Unseen split. When the maximum number of action chunks is equal to $1$, we set the merging probability to $1$ to make it equivalent to the action merging method used in previous works. We found that optimal performance was achieved when the model was allowed to output $3$ action chunks at a time, while fewer or more chunks resulted in degraded performance. We believe that appropriately increasing the number of action chunks predicted by the model can improve its long-term path planning capabilities, while predicting too many action chunks would exceed its perception limit, leading to suboptimal action predictions due to a lack of global knowledge.
Table 4: Ablation study of the maximum chunk number in HPAC on VLN-CE R2R Val-Unseen split.
@@
1 5.08 62.2 54.4 48.8
2 4.94 60.8 54.2 48.9
3 4.91 63.2 55.5 49.6
4 5.23 60.2 52.5 46.4
Entropy-Based vs. Fixed Execution Horizon. Tab. 5 presents our ablation study of entropy-based action execution horizon selection. As a comparison, we truncate the action chunks after a fixed number of action chunks. It can be seen that selecting an appropriate execution horizon (the second row) effectively improves the model’s performance in unseen scenarios. Furthermore, the action entropy-based execution horizon selection method is more effective than fixed-size execution horizon selection, demonstrating the benefit of adapting to the model’s uncertainty dynamics.
Table 5: Ablation study of the execution horizon on VLN-CE R2R Val-Unseen split.
[…]
Truncated@chunk 1 5.09 62.9 55.5 49.3
Truncated@chunk 2 4.93 63.3 56.2 50.4
Truncated@chunk 3 4.91 63.2 55.5 49.6
Entropy-based selection 4.70 64.5 57.4 51.2
Inverse vs. Forward Dynamics. To evaluate the effectiveness of inverse dynamics supervision, we compared it with forward dynamics supervision (FDS). Specifically, we follow MonoDream [34] and use a VLM to predict the visual features of the future without introducing any additional modules. For fairness, both methods were trained with the same amount of dynamics data for one epoch. As shown in Tab. 6, inverse dynamics supervision consistently outperforms forward dynamics supervision and the baseline model without any dynamic supervision, while forward dynamics supervision introduces significant performance degradation. We believe this difference stems from the consistency of the output space: enabling a pre-trained VLM to predict dense future visual features is typically harder to optimize and demands significant additional computation and data—for instance, MonoDream [34] requires training for five epochs. In contrast, inverse dynamics supervision aligns with the pre-training objective of the VLM, simplifying the learning of the causal relationship between visual changes and corresponding actions.
@@ 5 Conclusion
In this work, we presented NaVIDA, a unified VLN framework that integrates policy learning with action-grounded visual dynamics and adaptive execution. Through chunk-based inverse-dynamics supervision, NaVIDA learns the causal relationship between visual transitions and corresponding actions, enabling more interpretable and stable navigation. The proposed hierarchical probabilistic action chunking mechanism structures navigation trajectories into multi-level motion units, providing richer and longer-range supervision for inverse dynamics learning. During inference, an entropy-guided adaptive execution mechanism dynamically adjusts the action horizon to prevent error accumulation and enhance behavioral stability. Extensive experiments demonstrate that NaVIDA achieves superior navigation performance with fewer parameters compared to state-of-the-art methods, and real-world robot evaluations confirm its effectiveness and practical applicability.
[…]
Supplementary Material
6 Prompts for Different Tasks
@@
For the forward-dynamics supervision in Sec. 4.3, the requirement is to predict the visual observation features of the next time step based on the given historical frames and language instructions. We use the following prompt:
@@
7 More Implementation Details
7.1 Data Details
Fig. 6 depicts the distribution of atomic actions per HPAC chunk in our inverse-dynamics (IDM) dataset. The majority of chunks contain $3$ to $5$ atomic actions, consistent with the fixed chunk sizes used in prior works [46, 8, 39], while a smaller number of samples contain either fewer or more atomic actions. Longer chunks encode larger egocentric viewpoint shifts and parallax, yielding strong, informative supervision signals for inverse dynamics and supporting long-horizon planning by offering the model an extended temporal context. In contrast, shorter chunks emphasize subtle, fine-grained visual changes (e.g., small turns or local re-alignments), which facilitate precise control and reduce action ambiguity at short horizons. This natural mixture of long and short chunks serves as a complementary curriculum: HPAC contributes global structural cues for planning alongside local adjustment priors for stabilization, thereby strengthening the causal grounding between visual observations and actions and improving robustness to compounding errors during execution.
Figure 6: Distribution of atomic actions per HPAC chunk in the IDM data. Most chunks contain 3–5 atomic actions.
We also normalized the navigation process to a range of 0 to 1, and  Fig. 7 shows the distribution of action chunks as the navigation process progresses. It can be seen that in the early stages of navigation, HPAC merges a large number of atomic actions, which decreases slightly and stabilizes in the middle stages. In the later stages of navigation, closer to the target, HPAC tends to use fewer atomic actions to stabilize the robot’s position.
Figure 7: Distribution of actions on navigation progress.
7.2 Training Hyperparameters
NaVIDA is implemented on Qwen-2.5-VL-3B; the vision encoder is frozen and only the projector and LLM are trainable. We use AdamW with lr = 2e-5, BF16, Flash-Attention v2, global batch size = 64; additional hyperparameters are summarized in  Tab. 7.

### v2

@@ 4 Experiment
NaVILA [navila] RSS2025 8B $\checkmark$ 6.77 49.3 44.0 58.8
Uni-NaVid [uni-navid] RSS2025 7B $\checkmark$ 6.24 48.7 40.9 -
Aux-Think [aux-think] NeurIPS2025 8B $\checkmark$ 6.24 52.2 40.2 -
StreamVLN [streamvln] ICRA2026 7B $\checkmark$ 6.22 52.9 46.0 61.9
MonoDream [monodream] AAAI2026 2B $\checkmark$ 6.38 49.4 40.9 -
InternVLA-N1(S2) [internnav2025] - 7B $\checkmark$ 6.41 49.5 41.8 62.6
InternVLA-N1(S1+S2) [internnav2025] - 8B $\checkmark$ $\checkmark$ 5.91 53.5 46.1 65.3
NavFoM [navfom] ICLR2026 7B $\checkmark$ 5.51 57.4 49.4 60.2
JanusVLN [janusvln] ICLR2026 8B $\checkmark$ 6.06 56.2 47.5 62.1
@@ 4.1 Experiments Setup
Implementation Details. We implement NaVIDA on top of Qwen2.5-VL-3B [qwen2_5]. During training, the vision encoder is frozen, while the projector and LLM backbone are optimized with a learning rate of 2e52e-5 for one epoch over mixed VLN and inverse-dynamics data. Following NaVILA [navila], we cap the temporal context at eight historical frames; if a sequence exceeds this limit, we uniformly subsample it to eight frames to preserve coverage across the full trajectory. Unless otherwise stated, HPAC uses merge probability $p=0.7$ and a maximum level-2 chunk size of $n=3$. At inference, we only execute the first two predicted action chunks. Further hyperparameter details are provided in the Supplementary Material.
Evaluation Benchmarks. We evaluate NaVIDA on VLN-CE using the val-unseen splits of R2R-CE [r2r, vlnce] and RxR-CE [rxr, vlnce]. Following standard practice [janusvln, internnav2025, aux-think], we report Navigation Error (NE), Success Rate (SR), Oracle Success rate (OS), Success weighted by Path Length (SPL) and normalized Dynamic Time Warping (nDTW). These metrics jointly capture goal completion, trajectory efficiency, and path fidelity, and allow direct comparison to recent RGB-only and multi-sensor approaches. Detailed explanations of the metrics are provided in the Supplementary Material.
@@ 4.2 Main Results
Performance Evaluation. As shown in  Tab.˜1 (R2R-CE) and  Tab.˜2 (RxR-CE), NaVIDA achieves state-of-the-art results on the val-unseen splits using only a single RGB stream and a 3B VLM backbone. On R2R-CE, NaVIDA attains the best NE/OS/SR and the second-best SPL, surpassing the strongest 8B baseline JanusVLN [janusvln] by -0.46 NE, +4.3 OS, and +0.9 SR. On RxR-CE, NaVIDA matches the best SR (57.4) and achieves the best NE/SPL/nDTW, improving over StreamVLN and JanusVLN by +1.2 to +4.5 SR, +2.1 to +3.6 SPL, and -0.83 to -0.99 NE. The consistent gains on the longer and linguistically richer RxR-CE benchmark indicate stronger long-horizon alignment and reduced drift, enabled by IDS and hierarchical chunking. Compared with multi-sensor methods such as ETPNav and NavMorph, NaVIDA’s single-RGB policy remains competitive or superior on core metrics. It also achieves the best performance among RGB-only methods with only  35% of the parameters and without auxiliary generators or spatial encoders. Overall, these results demonstrate that inverse-dynamics augmentation with HPAC provides an effective and lightweight alternative to conventional map or forward world modeling.
Table 3: Ablation study of IDS Data on VLN-CE R2R Val-Unseen split.
@@
Table 4: Ablation study of the IDS Design Variants on VLN-CE R2R Val-Unseen split.
[…]
w/o IDS 5.29 61.9 52.6 47.1
MCQ 5.82 56.2 49.8 45.7
Noised IDS 6.21 53.9 46.8 42.7
PD 5.27 60.1 52.3 47.8
Ours 4.91 63.2 55.5 49.6
Qualitative Results. Fig.˜3 contrasts NaVIDA with a variant without IDS where green arrows mark correct actions and red arrows mark errors. NaVIDA correctly defers turning until clearing the staircase and selects the intended intersection, whereas the ablated model exhibits premature turns and junction misclassification. These are typical symptoms of error accumulation when the model lacks the understanding of action-grounded visual dynamics and fails to internalize the visual consequences of its actions.
Fig.˜4 presents examples of real-world trials. NaVIDA executes tight turns at narrow corners and passes smoothly through single-person doorways; even with pedestrians ahead, it resists interference and holds a reliable heading. Taken together, the simulator and real-world evidence identify IDS as the key component behind the observed stability and scene generalization, outperforming the variant without inverse-dynamics supervision.
@@ 4.3 Ablation Study
In this section, we conduct ablations on the effectiveness of IDS, the effectiveness of HPAC and dynamic modeling methods. To ensure efficiency, we conduct experiments on data without ScaleVLN and report the performance on R2R-CE benchmark. Here for convenience, we execute all predicted action chunks in this section.
Scalability of IDS Data.  Tab.˜4 varies whether R2R, RxR, and DAgger trajectories contribute IDS pairs ($\bullet$) or not ($\circ$). Incorporating IDS consistently improves performance across all metrics and data configurations. On R2R alone, adding IDS increases SR from 32.0 to 40.5. When extending to R2R+RxR, SR further rises to 47.7. Introducing DAgger without IDS yields an SR of 52.6, while enabling IDS supervision across all three sources achieves 55.5 SR, accompanied by a lower NE (4.91) and higher SPL (49.6). Notably, the gains are most pronounced in data-scarce settings and continue to accumulate as additional trajectory sources are introduced. This scaling behavior indicates that IDS provides a complementary training signal rather than merely amplifying imitation supervision. Overall, these results suggest that IDS injects a transferable, language-agnostic notion of action-grounded visual dynamics into the model. Crucially, this capability integrates seamlessly with existing rollout-based paradigms such as DAgger, yielding additive improvements without requiring changes to the underlying imitation framework.
Table 5: Ablation study of HPAC on VLN-CE R2R Val-Unseen split.
@@
1 Probabilistic 5.08 62.2 54.4 48.8
2 Probabilistic 4.94 60.8 54.2 48.9
3 Probabilistic 4.91 63.2 55.5 49.6
3 DINOv2-based 6.29 56.1 44.7 39.5
4 Probabilistic 5.23 60.2 52.5 46.4
@@
Ablations of IDS Design. To better understand how IDS facilitates action predicition, we conduct ablations on its construction strategy (Tab.˜4). First, we corrupt IDS by replacing ground-truth actions with random ones during data construction (Noised IDS), thereby breaking the alignment between actions and visual transitions while keeping the VLN data and training scale unchanged. This leads to a significant drop in navigation performance, as noisy supervision introduces inconsistent training signals for action tokens, which confirms that IDS primarily benefits the model by reinforcing fine-grained action-visual dynamics consistency, rather than merely increasing data size. We further reformulate IDS as a multiple-choice question (MCQ), where the model selects the correct action from four candidates instead of directly generating the action token. Although action–grounded visual dynamics is preserved, navigation performance slightly decreases.While the MCQ format allows the model to rely on simple comparison or elimination among candidate options, the generative IDS requires the agent to directly produce the action token based solely on the visual change. Sharing the same generative action space with VLN decoding is therefore important, as it ensures that the transition knowledge learned during IDS can be directly applied to action prediction during navigation, leading to more robust behavior. We also implement a privileged distillation (PD) variant. A teacher model is first trained with IDS supervision, and its knowledge is then distilled into a student on the VLN task by minimizing the KL divergence between their action distributions. Despite this additional supervision, PD performs similarly to the baseline without dynamic supervision and remains inferior to NAVIDA. We attribute this limited improvement to how the understanding of action-grounded visual dynamics is transferred. PD enforces alignment only at the output distribution level, encouraging the student to imitate the teacher’s final predictions. Such post-hoc supervision treats visual dynamics knowledge as a fixed action signal injected after training. In contrast, NAVIDA jointly optimizes IDS and VLN within a unified framework, allowing action–visual alignment to be continuously learned and reinforced through shared representations. This tighter coupling enables action-grounded visual dynamics to directly shape navigation reasoning, yielding more effective learning than distillation-based approaches.
Figure 5: IDS score over prediction steps for successful and failed trajectories. Successful episodes exhibit consistently higher and more stable IDS scores, while failed episodes show gradual degradation, indicating the importance of accurate action-state alignment for long-horizon navigation.
[…]
$\Delta_{p}=\|\hat{\mathbf{p}}-\mathbf{p}\|_{2},\quad\Delta_{\theta}=\min\left(\left|\hat{\theta}-\theta\right|,\,2\pi-\left|\hat{\theta}-\theta\right|\right).$ (3)
The IDS score is then defined as:
$\mathrm{IDS_{Score}}=\exp\left(-\frac{\Delta_{p}}{\sigma_{p}}\right)\cdot\exp\left(-\frac{\Delta_{\theta}}{\sigma_{\theta}}\right),$ (4)
where $\sigma_{p}$ and $\sigma_{\theta}$ control the sensitivity to positional and angular deviations, respectively. In our implementation, we set $\sigma_{p}=0.1$ and $\sigma_{\theta}=\pi/12$.
Fig.˜5 shows the evolution of IDS scores over prediction steps for successful and failed trajectories under different overall success rates. Several consistent trends can be observed. At early stages, the IDS scores of successful and failed episodes are similar, indicating that short-term action prediction is generally reliable in both cases. As navigation proceeds, however, the two curves gradually diverge: successful trajectories maintain higher and more stable IDS scores, whereas failed ones exhibit a steady decline. This widening gap suggests that the understanding and accurate prediction of action-induced visual transitions becomes increasingly important for long-horizon navigation. Furthermore, settings with higher overall success rates consistently present higher and more stable IDS scores. This strong association indicates that IDS accuracy reflects the model’s capability to understand the visual dynamics of its actions. When this understanding is precise, action sequences remain coherent over time, leading to improved navigation performance. In contrast, lower IDS scores in failed trajectories imply weaker understanding of the action–visual dynamics, where small deviations accumulate and ultimately result in failure.
Ablations of HPAC.  Tab.˜6 presents an ablation study of HPAC on the VLN-CE R2R val-unseen split. We first analyze the effect of the maximum number of action chunks. When the maximum is set to $1$ with merging probability fixed to $1$, HPAC reduces to the deterministic action-merging strategy adopted in prior work [navid, navila]. The best performance is achieved when allowing up to $3$ action chunks under our default probabilistic merging scheme, while using fewer or more chunks degrades results. This suggests that moderately extended chunks enhance longer-horizon planning, whereas overly long chunks may exceed the model’s effective perception range and lead to suboptimal decisions. Importantly, the probabilistic merging strategy introduces controlled randomness into chunk boundaries, preserving diversity of action chunks and improving robustness to minor execution fluctuations. We also evaluate a deterministic alternative based on DINOv2 [dinov2] similarity, where adjacent frames are grouped according to visual similarity. This variant performs worst overall. We attribute this to the fact that DINOv2 primarily captures high-level semantic similarity, whereas navigation depends on geometric and viewpoint changes. For instance, observing the same object from different viewpoints may yield high DINOv2 similarity despite significant spatial displacement. Such semantically driven and deterministic merging can obscure motion cues and reduce segmentation diversity, resulting in weaker action supervision and inferior performance.
Inverse vs. Forward Dynamics. To evaluate the effectiveness of inverse dynamics supervision, we compare it with forward dynamics supervision (FDS), as both aim to enhance action-grounded visual dynamics modeling. Following MonoDream [monodream], FDS predicts future visual features using the same VLM without introducing additional modules. Considering its slower convergence, we train FDS for three epochs to assess both effectiveness and efficiency relative to IDS. As shown in Tab.˜6, IDS consistently outperforms both FDS and the baseline without dynamic supervision, while FDS initially causes noticeable performance degradation. We attribute this gap to differences in output space alignment and optimization difficulty. Predicting dense future visual features requires modeling high-dimensional appearance variations, which is challenging for a pre-trained VLM and typically demands more training (e.g., MonoDream trains for five epochs). In contrast, IDS shares the same output space as action prediction in VLN, making it better aligned with the model’s pre-training and downstream objectives. Although FDS gradually improves with additional training and eventually surpasses the baseline after three epochs, it still lags behind IDS, which achieves superior performance after only one epoch. These results indicate that IDS provides more stable and data-efficient supervision, whereas FDS requires longer optimization to learn reliable future-state predictions.
@@ 5 Conclusion
In this work, we presented NaVIDA, a unified VLN framework that integrates policy learning with action-grounded visual dynamics modeling. Through chunk-based inverse-dynamics supervision, NaVIDA achieves more interpretable and stable navigation. The proposed hierarchical probabilistic action chunking mechanism structures navigation trajectories into multi-level motion units, providing richer and longer-range supervision for inverse dynamics learning. Extensive experiments demonstrate that NaVIDA achieves superior navigation performance with fewer parameters compared to state-of-the-art methods, and real-world robot evaluations confirm its effectiveness and practical applicability.
Appendix 0.A Prompts for Different Tasks
@@
For the multiple-choice question (MCQ) of IDS, the prompt is like follows:
[…]
For the forward-dynamics supervision, the requirement is to predict the visual observation features of the next time step based on the given historical frames and language instructions. We use the following prompt:
@@
Appendix 0.B More Implementation Details
0.B.1 Data Details
Fig.˜6(a) depicts the distribution of atomic actions per HPAC chunk in our inverse-dynamics (IDM) dataset. Although most chunks contain $3$ to $5$ atomic actions, which consistent with the fixed chunk sizes used in prior works [navid, navila, streamvln], a non-negligible portion of samples contain either fewer or more atomic actions. Longer chunks encode larger egocentric viewpoint shifts and parallax, yielding strong, informative supervision signals for inverse dynamics and supporting long-horizon planning by offering the model an extended temporal context. In contrast, shorter chunks emphasize subtle, fine-grained visual changes (e.g., small turns or local re-alignments), which facilitate precise control and reduce action ambiguity at short horizons. This natural mixture of long and short chunks serves as a complementary curriculum: HPAC contributes global structural cues for planning alongside local adjustment priors for stabilization, thereby improving the model’s understanding of action-grounded visual dynamics and actions and improving robustness to compounding errors during execution.
We also normalized the navigation process to a range of 0 to 1, and  Fig.˜6(b) shows the distribution of action chunks as the navigation process progresses. It can be seen that in the early stages of navigation, HPAC merges a large number of atomic actions, which decreases slightly and stabilizes in the middle stages. In the later stages of navigation, closer to the target, HPAC tends to use fewer atomic actions to stabilize the robot’s position.
(a) Distribution of atomic actions per HPAC chunk in the IDM data.
(b) Distribution of actions on navigation progress.
Figure 6: Atomic action statistics of HPAC in the dataset.
0.B.2 Training Hyperparameters
NaVIDA is implemented on Qwen-2.5-VL-3B; the vision encoder is frozen and only the projector and LLM are trainable. We use AdamW with lr = 2e-5, BF16, Flash-Attention v2, global batch size = 64; additional hyperparameters are summarized in  Tab.˜7.